OpenStudy (anonymous):

Find the rate of increase (with respect to r) in the surface area (S=4πr2) of a spherical balloon when: (A) r=2 inches → Rate of increase = (B) r=4 inches → Rate of increase = (C) r=6 inches → Rate of increase =

OpenStudy (anonymous):

rate of change = first derivative since we are findign the rate of change with respect to radius, we find \(\displaystyle {dS\over dr}\)

OpenStudy (anonymous):

what will be the derivative?

OpenStudy (anonymous):

S = 4pi*r^2 dS/dt = 4pi*2*r*dr/dt

OpenStudy (anonymous):

not with "t" just with "r"

OpenStudy (anonymous):

there is no variable "t" in the system. the question said .. with respect to radius (r)

OpenStudy (anonymous):

ok so what do i need to do?

OpenStudy (anonymous):

@electrokid

OpenStudy (anonymous):

\[ {dS\over dr}=4\pi(2r)=8\pi r \]

OpenStudy (anonymous):

this is what is asked

OpenStudy (anonymous):

ok so the answer is 8pi r

OpenStudy (anonymous):

yep. that is the rate of increase of S with respect to r

OpenStudy (anonymous):

ok, thank you so much

OpenStudy (anonymous):

yw